find two consecutive integers whose sum is 100
Answer by LilSkittleMd(119)

To find the answers for the consecutive integers problems you take x (the number) and x+1 (its consecutive number) and equal it to the sum
In this case your equation would be
x+(x+1)=100
x+x+1=100
2x=99
x=49.5
x+1=50.5
Your two consecutive integers are 49.5 and 50.5